OpenStudy (anonymous):

Find the values of x in this equation: x -15/x= 2 A. -7,3 B. -5,2 C. -7,5 D. -2,5 E. -3,5

pooja195 (pooja195):

Multiply by the reciprocal x-15/x=x/x-15 Once you do that you get this \[\huge~\rm~x−15=2x \] Can you solve for x and tell me what you get? :)

OpenStudy (anonymous):

15=x?

pooja195 (pooja195):

Hmm you seemed to have made an error can you show how you got that?

OpenStudy (anonymous):

I look the x and subractted it on both sides

pooja195 (pooja195):

ok so -15=3x

OpenStudy (texaschic101):

x - 15 = 2x.....try subtracting x from both sides...and don't forget the signs

pooja195 (pooja195):

Whats next?

OpenStudy (anonymous):

Divide 3x/-15

pooja195 (pooja195):

And what do you get?

OpenStudy (texaschic101):

wait..where are you getting the 3x

OpenStudy (anonymous):

-5, so the answer is B?

pooja195 (pooja195):

Wait 2-1-1x x−15=2x -15=1x divide from here sorry for the msitake

pooja195 (pooja195):

*mistake

OpenStudy (anonymous):

Thats not getting me an answer

pooja195 (pooja195):

-15/1=?

OpenStudy (anonymous):

-15

OpenStudy (texaschic101):

you could just try subbing in your answer choices to see which one is true

pooja195 (pooja195):

Good so x=-15

OpenStudy (anonymous):

I tried that nothing came out to 2

pooja195 (pooja195):

o.O it wants the values of x A. -7 -15/3= 2 Is this true or false?

OpenStudy (anonymous):

I guess its true

pooja195 (pooja195):

Dont guess try it out -7-15=?

OpenStudy (anonymous):

2

pooja195 (pooja195):

we are subtracting 15 from -7 -7 - 15 =?

OpenStudy (anonymous):

The problem might be confusing x - 15 OVER x = 2

pooja195 (pooja195):

I took option A (the values of x) And plugged them in for you

OpenStudy (texaschic101):

this makes no sense...your answer will never come out to be 2 because you are dividing a negative by a positive...resulting in a negative

OpenStudy (anonymous):

Well its a question sadly, I subbed them all in everyway i could and never got 2

pooja195 (pooja195):

wow...is this an online assignment? Is there a none of the above option?

OpenStudy (anonymous):

Yes, and yes

pooja195 (pooja195):

Then just hit none of the above :P

OpenStudy (anonymous):

Opps no for the second one

pooja195 (pooja195):

OH .-. yeah im not sure @texaschic101 is right though none are true

pooja195 (pooja195):

Did you mistype anything?

OpenStudy (anonymous):

Nope I hit E and went on

pooja195 (pooja195):

hmm sorry for improper help :/

OpenStudy (sleepyjess):

Is this the question? \(\dfrac{x-15}{x}=2\)

pooja195 (pooja195):

yea

pooja195 (pooja195):

Sorry we couldn't help you this time perhaps it was a mistake? \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

I have plenty more questions to lol

pooja195 (pooja195):

E is correct

pooja195 (pooja195):

x - (15/x) = 2 x^2 - 15 = 2x x^2 - 2x - 15 = 0 (x - 5)(x + 3) = 0 x = -3, 5

OpenStudy (texaschic101):

(x - 15) / x = 2/1 cross multiply 2x = x - 15 x = -15 I don't see any other answer it could be....oh....is it written like pooja wrote ?

pooja195 (pooja195):

It should have been written that way .-. that's the only possible answer

OpenStudy (sleepyjess):

no, she said it was (x-15)/x...

OpenStudy (sleepyjess):

that's the only way we can get an answer though

pooja195 (pooja195):

So congrats you got it correct! :D